OpenStudy (anonymous):

1. The distance traveled by an object can be modeled by the equation d = ut + 0.5at2 where d = distance, u = initial velocity, t = time, and a = acceleration. Solve this formula for a. Show all steps in your work.

OpenStudy (anonymous):

I found out the first step: Subtracting "ut" from both sides.

OpenStudy (anonymous):

@DebbieG

OpenStudy (zpupster):

so you have: d-ut= .5at2 the next step is to divide both sides by .5

OpenStudy (zpupster):

what do you get??

OpenStudy (anonymous):

Um I get d-ut/.5 = at2

ganeshie8 (ganeshie8):

still \(a\) is not free... isolate it completely

OpenStudy (anonymous):

You have to divide t2 to both sides

ganeshie8 (ganeshie8):

yup !

OpenStudy (anonymous):

But how do I do that since on the other side I have d - ut/.5???

OpenStudy (anonymous):

Wait, I looked this up and someone multiplied the other side by 2 instead of dividing by 0.5, I know how they did this. They changed 0.5 to a fraction (5/10) and simplified it to 1/2. So you can divide 1/2 to both sides or multiply 2/1(2) to both sides.

OpenStudy (anonymous):

I get it now.

ganeshie8 (ganeshie8):

cool :)

ganeshie8 (ganeshie8):

so wat do you get in the end

OpenStudy (anonymous):

a = 2(d – ut)/t2

OpenStudy (anonymous):

Is that correct?

ganeshie8 (ganeshie8):

looks perfect ! good job!!

OpenStudy (lena772):

When you multipy the other side by 2, does it look like 2(0.5at^2) ?

OpenStudy (lena772):

@ganeshie8 @greenlegodude57

undeadknight26 (undeadknight26):

@lena772 i do not believe so...

OpenStudy (anonymous):

Look you have d = ut + 0.5at2 First subtract "ut" from both sides: d - ut = 1/2at2. To get rid of the 1/2(0.5) you have to divide 1/2 to both sides or multiply 2/1(2) to both sides. So 2d - ut = at2. Now for your final step divide "t2" to both sides: 2d - ut/t2 = a. And you're done!